The Length Of A Rectangle Is 3 Times The Width The Perimeter Of The Field Is 144cm Find Its Area
Answer by ewatrrr(24785)

 
Hi,Previously Posted

P = 2L + 2w = 144cm
L+ w = 72 L = 3w
3w + w = 72
4w = 72
w = 18cm and L = 54cm
A = Lw = 972cm^2